### v1

# A Thermodynamic Theory of Learning I: Irreversible Ensemble Transport and Epistemic Costs

Abstract
Learning systems acquire structured internal representations through training, yet classical information-theoretic results assert that deterministic transformations cannot increase information. This apparent tension raises a fundamental question: how can learning produce meaningful epistemic structure without violating information-theoretic limits (e.g., data processing inequalities)?
In this work, we argue that this tension is resolved by recognizing learning as an inherently finite-time and irreversible process. We formulate learning at the ensemble level as a transport process in the space of probability distributions over model configurations. Within this framework, we introduce an epistemic free-energy functional that balances objective-driven potential improvement against the loss of ensemble diversity.
We show that reductions in epistemic free energy accumulated along a learning trajectory admit a decomposition into reversible contributions associated with changes in the objective landscape and irreversible contributions quantified by entropy production. Crucially, this entropy production reflects finite-time epistemic commitment and cannot be eliminated by algorithmic optimization alone.
Building on this formulation, we derive an Epistemic Speed Limit (ESL), a finite-time inequality that lower-bounds the irreversible entropy production required to realize a prescribed ensemble transformation. The bound is geometric in nature, depending only on the transport distance between initial and final distributions, and holds independently of the specific learning algorithm.
Our results establish a thermodynamic theory of learning in which differences between learning procedures arise not from access to different epistemic resources, but from how efficiently they manage unavoidable irreversible cost. This perspective reconciles abstraction with information-theoretic constraints and highlights finite-time irreversibility as a central organizing principle of learning dynamics.
footnotetext: This paper is the first part of a planned series entitled “A Thermodynamic Theory of Learning,” which develops a thermodynamic framework for finite-time learning dynamics.
1 Introduction
Learning systems acquire structured internal representations from data. In modern machine learning, this process enables models to extract abstractions, discover latent regularities, and generalize beyond observed examples. At the same time, classical results in information theory assert that deterministic transformations cannot increase information content. This apparent tension raises a fundamental question: how can learning give rise to meaningful structure and abstraction without contradicting established information-theoretic principles?
A key observation of this work is that learning is inherently a finite-time process: meaningful learning requires committing to epistemic structure within finite time Finzi et al. (2026). While asymptotic analyses often focus on the existence or properties of optimal solutions, real learning systems must operate under strict constraints on time, data, and computation. Under such constraints, learning dynamics are generally irreversible: probability mass is compressed, hypotheses are discarded, and epistemic uncertainty is reduced in ways that cannot be undone without additional cost. Understanding the nature and consequences of this irreversibility is therefore essential for a principled theory of learning.
In this work, we adopt a thermodynamic perspective on learning. We describe the state of a learning system by a probability distribution over model configurations and model learning as a transport process in the space of such distributions. Within this framework, we introduce an epistemic free-energy functional that balances an objective-driven contribution—captured by a learning objective such as empirical risk or negative log-likelihood—against epistemic uncertainty over model configurations. This formulation provides a unified language for quantifying both progress in learning and the costs incurred along a learning trajectory.
A central quantity in our analysis is the free-energy reduction, defined as a bookkeeping quantity that records the cumulative change in epistemic free energy along a learning trajectory. Importantly, this quantity specifies how much free energy is reduced, but not how that reduction is realized dynamically. To resolve this distinction, we analyze the decomposition of free-energy change into reversible and irreversible contributions, with the irreversible contribution corresponding to entropy production.
While the formalism developed here is grounded in thermodynamic concepts, it is important to clarify the role played by entropy production in learning. Entropy production should not be interpreted as directly determining the final performance of a learning system. Rather, it characterizes the irreversibility of the learning trajectory under finite-time constraints. In an idealized regime where learning dynamics can be approximated as gradient flows of a time-independent free-energy functional without external driving, the total free-energy reduction along a learning trajectory coincides with the total entropy production in this limit. This regime provides a useful baseline in which the irreversible cost of learning can be cleanly isolated.
Practical learning algorithms, however, rarely operate under such idealized conditions. Common training practices—including time-dependent learning-rate schedules, curriculum learning, and adaptive regularization— effectively introduce external driving that modifies the free-energy landscape over time. From this perspective, entropy production does not constrain which configurations a learning system may ultimately reach, but instead limits how efficiently a given distributional transformation can be realized within finite time. Excessive entropy production is therefore expected to manifest not as degraded asymptotic performance, but as instability, sensitivity to hyperparameters, or poor reproducibility of learning outcomes.
Building on this framework, we derive an Epistemic Speed Limit (ESL), a finite-time inequality that lower-bounds the entropy production required to transport probability mass between distributions over a given time interval. Crucially, this bound is geometric in nature: it depends only on the statistical distance between the initial and final distributions and the duration of the learning process, and holds independently of the specific learning algorithm or form of external driving, including optimally controlled dynamics. The ESL thus reveals a fundamental constraint on learning dynamics, clarifying the minimal irreversible cost associated with committing to epistemic structure under finite-time constraints.
2 Related Work
Epiplexity and structural information.
The concept of epiplexity was recently introduced as a measure of the amount of epistemic structure available to a computationally bounded observer Finzi et al. (2026). This notion builds on a long tradition of research on minimum description length (MDL) Rissanen (1978, 1989); Grünwald (2007) and resource-bounded variants of Kolmogorov complexity Li and Vitányi (2008), which characterize structure in terms of compressibility under computational constraints. In this line of work, epiplexity is fundamentally a static, in-principle quantity: it quantifies how much epistemic structure is present independently of how learning is actually carried out, and is typically defined through optimal descriptions or representations. In contrast, the present work is concerned not with the amount of structure that can in principle be extracted, but with the irreversible costs incurred when committing to such structure along finite-time learning trajectories. The present work should be viewed as complementary to epiplexity, addressing not availability but realization.
Free energy and variational learning.
Free-energy formulations play a central role in variational inference and Bayesian learning, where learning is interpreted as the minimization of a variational free-energy objective Neal and Hinton (1998); Wainwright and Jordan (2008); Bishop (2006). These approaches provide a continuous, distributional relaxation of discrete model selection and MDL-based objectives, allowing learning processes to be described as trajectories in probability space. The present work adopts this free-energy perspective not as a normative learning principle, but as a dynamical formalism for analyzing learning as a finite-time transport process. Here, free energy is not primarily interpreted as a measure of subjective belief or uncertainty, but as a representational tool that enables the decomposition of learning progress into objective-driven contributions and irreversible entropy production along learning trajectories.
Finite-time constraints and speed limits.
Fundamental limits on the rate of state transformations have been extensively studied in nonequilibrium thermodynamics, where speed limits relate achievable transformations to entropy production and geometric properties of state space Shiraishi et al. (2018); Ito (2024). A particularly clear class of results establishes universal lower bounds on dissipation in terms of geometric distances between states Ikeda et al. (2025). The present work can be viewed as an epistemic analogue of these thermodynamic speed limits, with probability distributions over model configurations playing the role of physical states. Closely related geometric ideas also arise in optimal transport theory, which provides natural metrics on spaces of probability distributions Villani (2008). The Epistemic Speed Limit derived here applies these finite-time irreversibility principles to learning dynamics, yielding bounds on the minimal irreversible cost required to transform distributions over model configurations within finite time.
Learning procedure dependence.
The strong dependence of learning outcomes on training procedures has long been recognized in practice. Curriculum learning Bengio et al. (2009) and knowledge distillation Hinton et al. (2015) are prominent examples in which structured or gradual training procedures lead to substantial improvements, even without additional data. Despite their empirical success, principled theoretical accounts of why learning procedures matter remain limited. From the perspective developed here, such methods can be interpreted as shaping learning trajectories so as to improve the efficiency of probability transport in distribution space, thereby managing the entropy production associated with finite-time learning. Importantly, entropy production itself is not inherently detrimental: in non-convex landscapes, stochastic fluctuations—and the associated entropy production— may be essential for exploration and for escaping unfavorable basins. The Epistemic Speed Limit instead highlights the cost of uncontrolled entropy production, arising from geometrically inefficient transport. Within this framework, advanced training strategies can be understood as allocating entropy production toward necessary exploration while limiting avoidable irreversibility.
3 Learning as an Irreversible Thermodynamic Process
In this section, we formulate learning as a finite-time irreversible process at the level of ensembles of model configurations. The role of thermodynamic language here is explicitly descriptive rather than prescriptive: the quantities introduced below serve as bookkeeping devices for analyzing ensemble-level learning dynamics. They do not define a learning objective, nor do they prescribe an optimization algorithm.
Our goal is to make precise in what sense learning is irreversible, what is lost along a learning trajectory, and how such losses can be quantified independently of any particular notion of optimality or generalization. These considerations will form the foundation for the Epistemic Speed Limit derived in Section 4.
3.1 Ensemble-Level Description of Learning
We describe learning at the ensemble level by a probability distribution $q_{s}(\theta)$ over a configuration space $\Theta$, indexed by a rescaled time parameter $s\in[0,1]$. An ensemble represents the set of model configurations that can be realized under variations in initialization, data ordering, stochasticity, or other training conditions.
This probabilistic description is not intended to represent subjective uncertainty or Bayesian belief. Rather, it provides a geometric representation of how learning redistributes accessible configurations across parameter space as training progresses.
Viewed in this way, learning is generically irreversible. As training proceeds, probability mass concentrates and regions of configuration space are abandoned, so that the initial ensemble cannot be recovered without incurring additional cost. This ensemble-level irreversibility motivates the transport-based formulation introduced in the following subsection.
3.2 Learning as Distributional Transport
We model learning as a continuous transport of probability distributions $q_{s}(\theta)$ over model parameters, indexed by a rescaled time $s\in[0,1]$. The evolution of $q_{s}$ is governed by the continuity equation
$\partial_{s}q_{s}+\nabla\cdot(q_{s}v_{s})=0,$ (1)
where $v_{s}(\theta)$ denotes the probability velocity field in parameter space.
As a geometric measure of irreversible ensemble transport, we define the epistemic entropy production rate [^1] Throughout this paper, the term “entropy production” is used in an epistemic and transport-theoretic sense. Specifically, it refers to the quadratic action associated with probability transport in Wasserstein space. This notion should be distinguished from the conventional thermodynamic entropy production defined in stochastic thermodynamics via heat exchange with an external bath. as
$\sigma_{s}:=\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta,\qquad\sigma_{s}\geq 0.$ (2)
The quantity $\Sigma_{0:1}$ will serve as a geometric measure of irreversible probability transport and forms the basic cost functional underlying the Epistemic Speed Limit derived below.
$\Sigma_{0:1}\;:=\;\int_{0}^{1}\sigma_{s}\,ds.$ (3)
This quantity coincides with the Benamou–Brenier action associated with the $2$-Wasserstein distance and measures the minimal irreversible cost required to transport probability mass between ensemble states. Throughout this paper, entropy production is understood in this epistemic, transport-theoretic sense.
3.3 Fokker–Planck Dynamics as a Representative Model
In this section, we focus on an idealized regime in which the effective noise temperature $T$ is treated as constant. This assumption allows a clean dissipation identity and highlights the geometric nature of irreversible ensemble transport. In practical learning algorithms such as SGD, the effective noise scale may vary over time due to learning-rate schedules or batch-size changes; such time-dependent driving generically introduces additional terms in the free-energy balance. However, the Epistemic Speed Limit derived here concerns the minimal geometric cost of probability transport and does not rely on the detailed temporal profile of $T$.
We use Fokker–Planck dynamics as a representative and analytically tractable model of irreversible ensemble learning. This choice is made for clarity rather than realism. The Fokker–Planck equation provides a canonical setting in which irreversibility, entropy production, and transport geometry can be computed explicitly.
Concretely, we consider a continuous-time approximation of stochastic gradient descent. Let $\theta_{s}$ evolve according to the Langevin equation
$d\theta_{s}=-\nabla\Phi(\theta_{s})\,ds+\sqrt{2T}\,dW_{s},$ (4)
where $\Phi(\theta)$ is a learning objective, $T>0$ sets the scale of stochasticity, and $W_{s}$ denotes standard Brownian motion.
The induced ensemble distribution $q_{s}(\theta)$ evolves according to the Fokker–Planck equation
$\partial_{s}q_{s}=\nabla\cdot\!\left(q_{s}\nabla\Phi(\theta)\right)+T\,\Delta q_{s}.$ (5)
This dynamics can be interpreted as the Wasserstein gradient flow of the epistemic free-energy functional introduced below Jordan et al. (1998); Ambrosio et al. (2008).
We emphasize that Gaussian noise should be understood as a convenient approximation rather than a literal model of stochastic gradient noise. The results derived below rely not on Gaussianity itself, but on the more general facts that learning induces irreversible ensemble transport and that such transport incurs a geometric cost over finite time.
3.4 Epistemic Free Energy and Entropy Production
To analyze ensemble-level learning dynamics, we introduce the epistemic free energy
$\mathcal{F}[q]:=\mathbb{E}_{q}[\Phi]-TH[q],\qquad H[q]:=-\int q(\theta)\log q(\theta)\,d\theta.$ (6)
The first term measures the ensemble-averaged objective value, while the second term penalizes concentration of probability mass. The free energy thus balances objective improvement against loss of ensemble diversity.
Under Fokker–Planck dynamics, the epistemic free energy $\mathcal{F}[q_{s}]$ satisfies a dissipation identity relating its rate of change to the entropy production rate $\sigma_{s}$ defined above:
$\frac{d}{ds}\mathcal{F}[q_{s}]=-\,T\,\sigma_{s},\qquad\sigma_{s}\geq 0,$ (7)
where $\sigma_{s}$ denotes the instantaneous entropy production rate. Integrating over $s\in[0,1]$ yields
$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:1},\qquad\Sigma_{0:1}:=\int_{0}^{1}\sigma_{s}\,ds.$ (8)
Equation (8) expresses a fundamental irreversibility. In this special dynamical setting, the decrease in free energy accumulated over the learning trajectory is exactly accounted for by the entropy production. More generally, however, entropy production and free-energy change are distinct quantities: the former depends on the entire learning path, while the latter depends only on the endpoints. This identity is dynamical and does not depend on any notion of optimality.
3.5 Free-Energy Decomposition
Independently of the dynamics, the free-energy difference between two ensemble states admits the algebraic decomposition
$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\bigl(\mathbb{E}_{q_{0}}[\Phi]-\mathbb{E}_{q_{1}}[\Phi]\bigr)+T\bigl(H[q_{1}]-H[q_{0}]\bigr).$ (9)
The first term represents the change in the ensemble-averaged objective, while the second term captures the loss of entropy along the learning trajectory.
This decomposition is purely definitional. In particular, it does not imply that objective improvement and entropy reduction are causally related. The role of the free energy is to provide a unified bookkeeping device in which both contributions appear on equal footing. Constraints on learning dynamics therefore apply to the total free-energy change, rather than to its individual components.
3.6 Gaussian Noise as a Representative Approximation
Gaussian noise is assumed throughout this work for analytical clarity, as it leads to a Fokker–Planck description with a quadratic entropy production rate and a natural Wasserstein-2 geometry. However, Gaussianity is not essential to the conceptual structure of our results. Empirical studies indicate that stochastic gradient noise is often heavy-tailed, in which case the ensemble dynamics are more appropriately described by Lévy-type or fractional diffusion processes. While such dynamics modify the precise form of the transport cost and the associated geometry, they preserve the fundamental features emphasized here: learning induces irreversible probability transport, entropy-like quantities decrease monotonically, and finite-time ensemble transformations incur unavoidable geometric costs. We therefore view the Fokker–Planck formulation as a representative model rather than a literal description of stochastic gradient dynamics.
3.7 Transport Geometry and Irreversible Cost
We now relate the entropy production to the geometry of probability transport.
Let $(q_{s},j_{s})$ be a learning trajectory satisfying the continuity equation
$\partial_{s}q_{s}+\nabla\cdot j_{s}=0,$ (10)
with fixed endpoints $q_{0}$ and $q_{1}$. The associated probability velocity field is defined by $j_{s}=q_{s}v_{s}$.
A fundamental geometric lower bound on the irreversible cost of such transport is provided by optimal transport theory. Specifically, the squared $2$-Wasserstein distance $W_{2}(q_{0},q_{1})^{2}$ defines a metric on the space of probability distributions over $\Theta$ with finite second moments and admits the variational representation
$W_{2}(q_{0},q_{1})^{2}=\inf_{(q_{s},v_{s})}\int_{0}^{1}\!\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta\,ds,$ (11)
where the infimum is taken over all paths satisfying the continuity equation with fixed endpoints. This quantity can be interpreted as the minimal quadratic cost required to transport probability mass from $q_{0}$ to $q_{1}$; see Villani (2008) for details.
Since the actual learning trajectory $(q_{s},v_{s})$ is one admissible path in this variational problem, the entropy production accumulated along learning is bounded from below by $W_{2}(q_{0},q_{1})^{2}$. This lower bound expresses a purely geometric and algorithm-independent constraint on finite-time learning dynamics.
4 Epistemic Speed Limit
Learning is represented as a trajectory in the space of probability distributions over model configurations, evolving from an initial ensemble $q_{0}$ to a terminal ensemble $q_{1}$ over a finite time horizon. As discussed in the previous section, finite-time learning dynamics necessarily incur irreversible costs associated with probability transport.
In this section, we formalize this constraint for a specific and analytically tractable class of learning dynamics. Our result is an epistemic analogue of thermodynamic speed limits, which bound dissipation in terms of geometric distances between states in nonequilibrium systems Ikeda et al. (2025).
Specifically, we derive a finite-time lower bound on the entropy production required to realize a prescribed ensemble transformation under Fokker–Planck dynamics. We refer to this bound as the Epistemic Speed Limit (ESL).
Theorem 1 (Epistemic Speed Limit under Fokker–Planck Dynamics).
Let $q_{s}(\theta)$ evolve according to the Fokker–Planck equation
$\partial_{s}q_{s}=\nabla\cdot\!\left(q_{s}\nabla\Phi\right)+T\Delta q_{s},\qquad s\in[0,1],$
from an initial distribution $q_{0}$ to a terminal distribution $q_{1}$.
Let $\Sigma_{0:1}$ denote the entropy production accumulated along the learning trajectory, as defined in Section 3. Then the following statements hold:
  1. 1.
    (Free-energy dissipation)
    $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:1}.$
  2. 2.
    (Epistemic Speed Limit)
    $T\,\Sigma_{0:1}\;\geq\;W_{2}(q_{0},q_{1})^{2}.$
  3. 3.
    (Consequence for objective improvement)
    $\mathbb{E}_{q_{0}}[\Phi]-\mathbb{E}_{q_{1}}[\Phi]\;\geq\;W_{2}(q_{0},q_{1})^{2}+T\bigl(H[q_{0}]-H[q_{1}]\bigr).$
Moreover, the bound in (2) is tight: equality is achieved when the ensemble trajectory follows a constant-speed Wasserstein geodesic.
Proof.
The free-energy dissipation identity follows from the standard calculation for the Wasserstein gradient flow of $\mathcal{F}$.
For the speed limit, recall that the squared Wasserstein distance $W_{2}(q_{0},q_{1})^{2}$ is defined as the minimum quadratic transport cost over all admissible probability flows connecting $q_{0}$ and $q_{1}$. Since the actual learning trajectory $(q_{s},v_{s})$ is one such admissible flow, the entropy production accumulated along learning must satisfy
$T\,\Sigma_{0:1}=\int_{0}^{1}\!\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta\,ds\;\geq\;W_{2}(q_{0},q_{1})^{2}.$
The final inequality follows directly from the algebraic decomposition of the free energy. ∎
Remark 1 (Finite-time scaling of the Epistemic Speed Limit).
Throughout this work, learning trajectories are parametrized by a normalized time variable $s\in[0,1]$. To relate the Epistemic Speed Limit to a physical training time $\tau\in[0,\mathcal{T}]$, consider the time reparametrization $\tau=\mathcal{T}s$. Under this change of variables, the continuity equation remains invariant, while the probability velocity field rescales as $v_{\tau}=(1/\mathcal{T})\,v_{s}$.
As a consequence, the entropy production accumulated over physical time satisfies
$\Sigma_{0:\mathcal{T}}=\frac{1}{T}\int_{0}^{\mathcal{T}}\!\int q_{\tau}(\theta)\,\|v_{\tau}(\theta)\|^{2}\,d\theta\,d\tau=\frac{1}{\mathcal{T}}\,\Sigma_{0:1}.$
This inequality shows that, for a fixed ensemble transformation, the minimal irreversible cost diverges as the available physical training time $\mathcal{T}$ decreases.
Combining the free-energy dissipation identity $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:\mathcal{T}}$ with the physical-time ESL yields the corresponding bound on free-energy drop:
$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]\;\geq\;\frac{1}{\mathcal{T}}\,W_{2}(q_{0},q_{1})^{2}.$ (12)
5 Discussion
The Epistemic Speed Limit (ESL) provides a unified perspective on learning efficiency by reframing learning not in terms of attainable outcomes, but in terms of the irreversible epistemic cost required to realize them within finite time. By separating epiplexity as an available epistemic resource from the entropy production incurred during learning, the ESL establishes that any finite-time learning process necessarily incurs irreversible epistemic entropy production. As a consequence, learning outcomes can differ across learning procedures—not because they access different information, but because they differ in how efficiently they convert available epistemic structure into useful potential improvement.
5.1 Learning Efficiency and Irreversible Cost
From the perspective of the ESL, learning corresponds to a trajectory in the space of probability distributions over model configurations. Different learning procedures induce different trajectories, which may incur substantially different amounts of irreversible entropy production even when they achieve comparable final performance.
Learning procedures that follow geometrically inefficient trajectories incur large entropy production in stabilizing internal representations. In contrast, procedures that follow smoother or more direct trajectories can realize comparable learning outcomes at lower irreversible cost. Differences in learning efficiency therefore reflect differences in entropy production rather than differences in accessible information or epistemic resources.
While the Epistemic Speed Limit constrains the irreversible cost of learning trajectories, improved transport efficiency does not necessarily imply improved generalization performance. In some regimes, solutions that are far from initialization may incur higher transport cost yet exhibit superior generalization. The ESL therefore constrains the cost of realizing ensemble transformations, rather than prescribing which endpoints are desirable from a performance perspective.
Importantly, the ESL does not propose new learning algorithms. Rather, it provides a principled explanation for why improvements in learning often arise from reducing unnecessary irreversible cost, rather than from accessing additional information.
5.2 Curriculum Learning, Distillation, and Teacher Guidance
Curriculum learning, knowledge distillation, and teacher guidance admit a natural interpretation within the ESL framework. These methods do not increase epiplexity or introduce new information. Instead, they reshape learning trajectories so as to avoid unnecessary irreversible entropy production.
Curriculum learning guides the learner along smoother paths in distribution space, reducing abrupt shifts that would otherwise require large entropy production. Similarly, knowledge distillation and teacher guidance constrain the learning trajectory, preserving epistemic flexibility and reducing entropy production during learning. Their empirical effectiveness follows from their ability to guide learning closer to quasi-reversible trajectories.
Learning Speed and Quasi-Static Limits
The finite-time form of the Epistemic Speed Limit also clarifies the role of learning speed. As the available training time $\mathcal{T}$ increases, the minimal entropy production required to realize a given ensemble transformation decreases, vanishing in the quasi-static limit $\mathcal{T}\to\infty$. In this idealized regime, learning can in principle proceed in an approximately reversible manner.
The ESL therefore does not prohibit low-dissipation learning. Rather, it quantifies the fundamental trade-off between learning speed and irreversible cost under finite-time constraints. In practical settings, where training time, computational resources, and environmental stability are limited, learning necessarily incurs nonzero entropy production. Excess dissipation then reflects geometrically inefficient learning trajectories rather than unavoidable thermodynamic cost.
5.3 Continual Learning and Irreversible Commitment
The perspective offered by the Epistemic Speed Limit also clarifies why subsequent learning after convergence is often difficult. Once learning has progressed to a terminal ensemble, the distribution over model configurations is typically highly concentrated, corresponding to a low-entropy state. This concentration reflects an irreversible epistemic commitment: probability mass has been transported away from large regions of configuration space, and those regions become effectively inaccessible under finite-time constraints.
Importantly, what is lost through learning is not information in the information-theoretic sense, but reachability. A low-entropy ensemble does not imply that alternative solutions do not exist in principle, but rather that reaching them from the current ensemble requires large probability transport. Under the Epistemic Speed Limit, such transport necessarily incurs substantial irreversible entropy production when attempted within finite time.
From this viewpoint, the difficulty of re-learning or adapting to new objectives is not primarily a consequence of insufficient data or model capacity. Instead, it arises because the ensemble has already undergone a geometrically costly redistribution. If the configurations that would reduce a new objective lie far outside the support of the current ensemble, the Wasserstein distance between the current and desired distributions is large, and rapid adaptation becomes intrinsically expensive.
Conversely, ensembles with higher entropy retain broader support in parameter space, making them closer, in a transport-geometric sense, to a wider range of potential future objectives. Such ensembles admit lower-cost adaptation, not because they encode more information, but because they preserve reachability. This observation highlights that entropy should not be viewed as something to be maximized or minimized in isolation. Rather, entropy mediates a trade-off between objective-driven improvement and future adaptability.
The Epistemic Speed Limit thus reframes the stability–plasticity dilemma as a geometric trade-off. Excessive concentration leads to efficient short-term optimization but incurs large transport costs for subsequent learning, while excessive dispersion slows objective-driven progress. Effective learning procedures implicitly balance these competing demands by shaping learning trajectories that reduce unnecessary irreversible cost while maintaining sufficient distributional flexibility for future adaptation.
5.4 Implications for Intelligence Growth
The following discussion should be understood as a conceptual implication rather than a concrete bound on any specific learning system. Here, “intelligence growth” refers abstractly to the accumulation of epistemic structure over finite time, rather than to any particular operational metric or task performance.
Beyond specific learning procedures, the ESL implies a fundamental constraint on the growth of intelligence over finite time. Even if epiplexity is unbounded in principle—through increasing data, model capacity, or representational richness—the realization of epistemic structure necessarily incurs irreversible entropy production.
This does not preclude rapid improvements in learning performance. Rather, it constrains the efficiency with which such improvements can be realized. Claims of instantaneous or arbitrarily explosive intelligence growth implicitly assume vanishing entropy production, infinite learning time, or a breakdown of the assumptions underlying irreversible learning dynamics.
Viewed through this lens, intelligence is shaped not only by the availability of epistemic structure, but by the history of irreversible epistemic commitments made during learning. The Epistemic Speed Limit thus highlights finite-time efficiency and irreversibility as central organizing principles of learning and intelligence.
6 Conclusion
In this work, we have argued that learning should be understood not only in terms of what epistemic structure is available in principle, but in terms of the irreversible cost required to realize such structure through finite-time dynamics. While epiplexity characterizes the epistemic structure accessible under computational and representational constraints, actual learning necessarily involves irreversible commitment to specific model configurations, accompanied by unavoidable entropy production.
To make this distinction explicit, we introduced a free-energy formulation in which learning is represented as a transport process over ensembles of model configurations. Within this framework, reductions in free energy can be algebraically decomposed into changes in the ensemble-averaged objective and changes in entropy. Irreversible entropy production, by contrast, is a path-dependent quantity that accounts for the free-energy reduction realized along finite-time learning trajectories. This decomposition highlights a central fact: meaningful learning over finite time cannot be realized without incurring an irreversible epistemic cost.
Our main theoretical result, the Epistemic Speed Limit (ESL), formalizes this observation as a finite-time law. The ESL does not bound learning outcomes or restrict what can be learned in principle. Instead, it lower-bounds the irreversible entropy production required to realize a prescribed ensemble transformation within a given time horizon. This bound is geometric in nature, determined by the transport cost between initial and final ensemble distributions, and holds independently of the specific learning algorithm employed.
The Epistemic Speed Limit also clarifies why further learning after apparent convergence is often intrinsically difficult. Once a learning process has completed, the ensemble over model configurations is typically highly concentrated, corresponding to a low-entropy state. This concentration reflects an irreversible epistemic commitment: probability mass has been transported away from large regions of configuration space, which thereafter become effectively unreachable under finite-time constraints.
Crucially, what is lost through learning is not information in the information-theoretic sense, but reachability. Alternative configurations that would support different objectives may still exist in principle, yet adapting to them from a concentrated ensemble requires large distributional transport. Under the Epistemic Speed Limit, such transport necessarily incurs substantial irreversible cost. From this perspective, the difficulty of continual learning and re-adaptation arises not from insufficient data or model capacity, but from the irreversible history of epistemic commitments accumulated during prior learning.
From this perspective, differences between learning procedures arise not because they access different epistemic resources, but because they manage unavoidable irreversible cost with differing efficiency. Procedures such as curriculum learning, knowledge distillation, and teacher guidance do not circumvent information-theoretic limits. Rather, they reshape learning trajectories so as to reduce unnecessary entropy production, allowing a larger fraction of available epistemic structure to be converted into realized potential improvement.
More broadly, the ESL implies a fundamental constraint on learning and intelligence over finite time. Even if epistemic structure is abundant or unbounded in principle, its realization necessarily incurs irreversible entropy production. Claims of instantaneous or arbitrarily explosive intelligence growth therefore implicitly assume vanishing dissipation, infinite learning time, or a breakdown of the assumptions underlying irreversible learning dynamics.
Viewed in this way, learning and intelligence are governed not solely by the availability of information or representational capacity, but by finite-time efficiency constraints imposed by irreversible epistemic commitment. The Epistemic Speed Limit thus identifies entropy production and transport geometry as central organizing principles of learning dynamics, complementing classical information-theoretic accounts with a fundamentally dynamical perspective.

### v2

@@
 where $\Phi(\theta)$ is a learning objective, $T>0$ sets the scale of stochasticity, and $W_{s}$ denotes standard Brownian motion.
-The induced ensemble distribution $q_{s}(\theta)$ evolves according to the Fokker–Planck equation
+This dynamics can be written in continuity form as
+$\partial_{s}q_{s}+\nabla\cdot(q_{s}v_{s})=0,$ (5)
+with the associated probability-flow velocity field
+$v_{s}=-\nabla\bigl(\Phi+T\log q_{s}\bigr).$ (6)
+Substituting this expression for $v_{s}$ into the continuity equation recovers the Fokker–Planck equation below.
 $\partial_{s}q_{s}=\nabla\cdot\!\left(q_{s}\nabla\Phi(\theta)\right)+T\,\Delta q_{s}.$ (7)
@@
 Under Fokker–Planck dynamics, the epistemic free energy $\mathcal{F}[q_{s}]$ satisfies a dissipation identity relating its rate of change to the entropy production rate $\sigma_{s}$ defined above:
-$\frac{d}{ds}\mathcal{F}[q_{s}]=-\,T\,\sigma_{s},\qquad\sigma_{s}\geq 0,$ (7)
-where $\sigma_{s}$ denotes the instantaneous entropy production rate. Integrating over $s\in[0,1]$ yields
-$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:1},\qquad\Sigma_{0:1}:=\int_{0}^{1}\sigma_{s}\,ds.$ (8)
+$\frac{d}{ds}\mathcal{F}[q_{s}]=-\sigma_{s},\qquad\sigma_{s}\geq 0,$ (9)
+where $\sigma_{s}$ denotes the instantaneous entropy production rate.
+A complete derivation of the dissipation identity, including all regularity assumptions and integration-by-parts steps, is provided in Appendix A.
+Integrating over $s\in[0,1]$ yields
+$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\Sigma_{0:1},\qquad\Sigma_{0:1}:=\int_{0}^{1}\sigma_{s}\,ds.$ (10)
 Equation (10) expresses a fundamental irreversibility.
@@
 $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\bigl(\mathbb{E}_{q_{0}}[\Phi]-\mathbb{E}_{q_{1}}[\Phi]\bigr)+T\bigl(H[q_{1}]-H[q_{0}]\bigr).$ (11)
-The first term represents the change in the ensemble-averaged objective, while the second term captures the loss of entropy along the learning trajectory.
+The first term represents the change in the ensemble-averaged objective, while the second term accounts for the change in entropy along the learning trajectory (it is negative when entropy decreases).
 This decomposition is purely definitional.
@@
 Empirical studies indicate that stochastic gradient noise is often heavy-tailed, in which case the ensemble dynamics are more appropriately described by Lévy-type or fractional diffusion processes.
-While such dynamics modify the precise form of the transport cost and the associated geometry, they preserve the fundamental features emphasized here: learning induces irreversible probability transport, entropy-like quantities decrease monotonically, and finite-time ensemble transformations incur unavoidable geometric costs.
+While such dynamics modify the precise form of the transport cost and the associated geometry, they preserve the fundamental features emphasized here: learning induces irreversible probability transport, the free energy decreases monotonically, and finite-time ensemble transformations incur unavoidable geometric costs.
 We therefore view the Fokker–Planck formulation as a representative model rather than a literal description of stochastic gradient dynamics.
@@
 from an initial distribution $q_{0}$ to a terminal distribution $q_{1}$.
-Let $\Sigma_{0:1}$ denote the entropy production accumulated along the learning trajectory, as defined in Section 3. Then the following statements hold:
+Let $\Sigma_{0:1}:=\int_{0}^{1}\sigma_{s}\,ds$ denote the entropy production accumulated along the learning trajectory, where $\sigma_{s}$ is defined in Section 3. Then the following statements hold:
   1. 1.
     (Free-energy dissipation)
-    $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:1}.$
+    $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\Sigma_{0:1}.$ (14)
   2. 2.
     (Epistemic Speed Limit)
-    $T\,\Sigma_{0:1}\;\geq\;W_{2}(q_{0},q_{1})^{2}.$
+    $\Sigma_{0:1}\;\geq\;W_{2}(q_{0},q_{1})^{2}.$ (15)
   3. 3.
     (Consequence for objective improvement)
-    $\mathbb{E}_{q_{0}}[\Phi]-\mathbb{E}_{q_{1}}[\Phi]\;\geq\;W_{2}(q_{0},q_{1})^{2}+T\bigl(H[q_{0}]-H[q_{1}]\bigr).$
-Moreover, the bound in (2) is tight: equality is achieved when the ensemble trajectory follows a constant-speed Wasserstein geodesic.
+    $\mathbb{E}_{q_{0}}[\Phi]-\mathbb{E}_{q_{1}}[\Phi]\;\geq\;W_{2}(q_{0},q_{1})^{2}+T\bigl(H[q_{0}]-H[q_{1}]\bigr).$ (16)
+Moreover, the bound in (15) is tight: equality is achieved for a constant-speed Wasserstein geodesic between $q_{0}$ and $q_{1}$, although such a trajectory does not necessarily coincide with the Fokker–Planck dynamics for a fixed objective $\Phi$.
 Proof.
@@
 Since the actual learning trajectory $(q_{s},v_{s})$ is one such admissible flow, the entropy production accumulated along learning must satisfy
-$T\,\Sigma_{0:1}=\int_{0}^{1}\!\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta\,ds\;\geq\;W_{2}(q_{0},q_{1})^{2}.$
-The final inequality follows directly from the algebraic decomposition of the free energy. ∎
+$\Sigma_{0:1}=\int_{0}^{1}\!\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta\,ds\;\geq\;W_{2}(q_{0},q_{1})^{2}.$
+The final inequality follows by combining the Epistemic Speed Limit with the algebraic decomposition of the free energy. ∎
 Remark 1 (Finite-time scaling of the Epistemic Speed Limit).
@@
 As a consequence, the entropy production accumulated over physical time satisfies
-$\Sigma_{0:\mathcal{T}}=\frac{1}{T}\int_{0}^{\mathcal{T}}\!\int q_{\tau}(\theta)\,\|v_{\tau}(\theta)\|^{2}\,d\theta\,d\tau=\frac{1}{\mathcal{T}}\,\Sigma_{0:1}.$
-This inequality shows that, for a fixed ensemble transformation, the minimal irreversible cost diverges as the available physical training time $\mathcal{T}$ decreases.
-Combining the free-energy dissipation identity $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:\mathcal{T}}$ with the physical-time ESL yields the corresponding bound on free-energy drop:
+$\Sigma_{0:\mathcal{T}}=\int_{0}^{\mathcal{T}}\!\int q_{\tau}(\theta)\,\|v_{\tau}(\theta)\|^{2}\,d\theta\,d\tau=\frac{1}{\mathcal{T}}\,\Sigma_{0:1}.$
+This scaling shows that, for a fixed ensemble transformation, the minimal irreversible cost diverges as the available physical training time $\mathcal{T}$ decreases.
+Combining this scaling with the free-energy dissipation identity $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\Sigma_{0:\mathcal{T}}$ yields the corresponding finite-time bound on free-energy reduction:
 $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]\;\geq\;\frac{1}{\mathcal{T}}\,W_{2}(q_{0},q_{1})^{2}.$ (17)
+This inequality should be interpreted as a geometric lower bound determined by the available learning time.
 5 Discussion
@@
 The Epistemic Speed Limit thus identifies entropy production and transport geometry as central organizing principles of learning dynamics, complementing classical information-theoretic accounts with a fundamentally dynamical perspective.
+Appendix A Monotonicity of the Epistemic Free Energy
+In this appendix, we provide a complete derivation showing that the epistemic free-energy functional decreases monotonically along the Fokker–Planck learning dynamics and that its rate of decrease coincides with the epistemic entropy production rate defined in the main text.
+Lemma 1 (Free-energy dissipation along Fokker–Planck dynamics).
+Let $q_{s}(\theta)$ be a smooth probability density evolving on $s\in[0,1]$ according to the Fokker–Planck equation
+$\partial_{s}q_{s}=\nabla\cdot\!\left(q_{s}\nabla\Phi(\theta)\right)+T\,\Delta q_{s},\qquad T>0,$ (18)
+where $\Phi(\theta)$ is a time-independent objective function. Define the epistemic free-energy functional
+$\mathcal{F}[q]:=\int q(\theta)\,\Phi(\theta)\,d\theta-T\,H[q],\qquad H[q]:=-\int q(\theta)\log q(\theta)\,d\theta.$ (19)
+Assume that $q_{s}$ is strictly positive and decays sufficiently fast at infinity (or satisfies appropriate boundary conditions) so that all integrations by parts are justified.
+Then the free energy satisfies
+$\frac{d}{ds}\mathcal{F}[q_{s}]=-\sigma_{s}\leq 0,$ (20)
+where
+$\sigma_{s}:=\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta,\qquad v_{s}(\theta):=-\nabla\!\left(\Phi(\theta)+T\log q_{s}(\theta)\right),$ (21)
+is the epistemic entropy production rate. Consequently,
+$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\int_{0}^{1}\sigma_{s}\,ds.$ (22)
+Proof.
+We differentiate the free-energy functional along the trajectory $q_{s}$. Using $\frac{d}{ds}\int q\Phi=\int\Phi\,\partial_{s}q$ and $\frac{d}{ds}\int q\log q=\int(1+\log q)\,\partial_{s}q$, we obtain
+$\frac{d}{ds}\mathcal{F}[q_{s}]=\int\bigl(\Phi+T(1+\log q_{s})\bigr)\,\partial_{s}q_{s}\,d\theta.$ (23)
+Substituting the Fokker–Planck equation (18) yields
+$\frac{d}{ds}\mathcal{F}[q_{s}]$ $=\int\bigl(\Phi+T(1+\log q_{s})\bigr)\Bigl[\nabla\cdot(q_{s}\nabla\Phi)+T\,\Delta q_{s}\Bigr]\,d\theta.$ (24)
+We treat the two terms separately. By integration by parts and the assumed boundary conditions,
+$\int\bigl(\Phi+T(1+\log q_{s})\bigr)\,\nabla\cdot(q_{s}\nabla\Phi)\,d\theta$ $=-\int q_{s}\,(\nabla\Phi+T\nabla\log q_{s})\cdot\nabla\Phi\,d\theta,$ (25)
+$T\int\bigl(\Phi+T(1+\log q_{s})\bigr)\,\Delta q_{s}\,d\theta$ $=-T\int(\nabla\Phi+T\nabla\log q_{s})\cdot\nabla q_{s}\,d\theta.$ (26)
+Using $\nabla q_{s}=q_{s}\nabla\log q_{s}$ and combining the two expressions, we obtain
+$\frac{d}{ds}\mathcal{F}[q_{s}]=-\int q_{s}(\theta)\,\|\nabla\Phi(\theta)+T\nabla\log q_{s}(\theta)\|^{2}\,d\theta.$ (27)
+Introducing the velocity field $v_{s}(\theta):=-\nabla(\Phi(\theta)+T\log q_{s}(\theta))$, this expression reduces to
+$\frac{d}{ds}\mathcal{F}[q_{s}]=-\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta=-\sigma_{s},$ (28)
+which proves (20). Integrating over $s\in[0,1]$ yields (22). ∎